OpenStudy (anonymous):

what is the simplified form of the expression (2x^6)(3x^1/2) my choices are: A.6x^13/2 B.6x^3 C.5x^13/2 D.5x^3 and thank you so much for helping btw, i suck at math :(

zepdrix (zepdrix):

So we multiply the 2 and the 3 together, which should eliminate a couple options for us. Then to multiply the x's we'll use this rule:\[\Large\bf\sf x^a\cdot x^b\quad=\quad x^{a+b}\]

zepdrix (zepdrix):

\[\Large\bf\sf (2x^6)(3x^{1/2})\quad=\quad 2\cdot3\cdot x^6\cdot x^{1/2}\quad=\quad 6x^{6+1/2}\]

zepdrix (zepdrix):

Understand how that rule works? We need to simplify the exponent now. To add the terms, you need a common denominator.